OpenStudy (anonymous):

the slope of a line is -1/3. The point (-12, -6) is on the line. Determine the equation of this line

OpenStudy (anonymous):

y=mx+b y=-6 x=-12 m=-1/3 -6=(-1/3*-12)+b -6=4+b b=-10 y=-1/3x-10